### v1

# NeurIPS 2025 Workshop on Space in Vision, Language, and Embodied AI

Abstract
Common computer vision systems typically assume ideal pinhole cameras but fail when facing real-world camera effects such as fisheye distortion and rolling shutter, mainly due to the lack of learning from training data with camera effects. Existing data generation approaches suffer from either high costs, sim-to-real gaps or fail to accurately model camera effects. To address this bottleneck, we propose 4D Gaussian Ray Tracing (4D-GRT), a novel two-stage pipeline that combines 4D Gaussian Splatting with physically-based ray tracing for camera effect simulation. Given multi-view videos, 4D-GRT first reconstructs dynamic scenes, then applies ray tracing to generate videos with controllable, physically accurate camera effects. 4D-GRT achieves the fastest rendering speed while performing better or comparable rendering quality compared to existing baselines. Additionally, we construct eight synthetic dynamic scenes in indoor environments across four camera effects as a benchmark to evaluate generated videos with camera effects.
Figure 1: We propose 4D Gaussian Ray Tracing (4D-GRT), a novel framework for generating physically-accurate, controllable camera effects in dynamic scenes. (1) Given multi-view video input, our method reconstructs a dynamic scene using 4D Gaussian Splatting (4D-GS) and differentiable ray tracing. (2) We simulate various camera effects with controllable parameters using ray tracing, generating high-quality videos with controllable camera effects.
Figure 2: We evaluate several state-of-the-art video generation models by specifying camera parameters in prompts to generate videos with specific effects. The results show that these models fail to generate physically accurate videos, instead producing artifacts or incorrect effects. Please refer to the supplementary materials for details.
1 Introduction
In the real world, various camera effects are usual phenomena that fundamentally characterize how cameras serve as the visual interface between the physical world and the digital world. For example, fisheye cameras capture substantially more scene content than conventional perspective cameras, providing additional context across a wider range of viewing directions.
However, common computer vision systems only consider pinhole cameras, rather than leveraging the beneficial properties of camera effects, because of the lack of high-quality visual data paired with accurate camera-effect parameters. As a result, these systems could fail when tested with these common and unavoidable camera effects in real-world scenarios. Gong et al. (2024); Fang et al. (2024); Fang and Chen (2025); Wang et al. (2023b) This data scarcity issue becomes a significant obstacle for modeling these effects, particularly in dynamic scenes where they interact non-trivially with motion (e.g., rolling shutter depends on object and scene dynamics Zhong et al. (2021); Saurer et al. (2013)). To overcome this issue, a straightforward approach is to synthesize data with camera effects. Traditional methods rely on physics-based rendering engines such as Blender Blender Online Community (2025), but they are labor intensive, costly, and can suffer from a pronounced sim-to-real gap Nikolenko et al. (2021); Tremblay et al. (2018). More recently, world models that are largely built on pretrained video generation models Ball et al. (2025); He et al. (2025); Hu et al. (2023, 2024) have been promoted as a path to realistic, controllable data. Unfortunately, as shown in Fig. 2, off-the-shelf video generation models have limited understanding of camera effects and their corresponding parameters, often generating videos that violate physical principles. These findings suggest that such world models underfit visual data paired with accurate camera-effect parameters, creating a chicken-and-egg problem.
The failure of these approaches motivates us to generate such video data by direct 4D reconstruction. With a reconstructed 4D representation, we can eliminate both the high cost of real-world data collection and the sim-to-real gap. However, existing reconstruction-based methods present significant shortcomings for this task. While Dynamic Gaussian Splatting methods Luiten et al. (2024); Yang et al. (2024b); Wu et al. (2024a) excel at dynamic scene reconstruction, they lack the capability to simulate light transport necessary for camera effect modeling. Conversely, Dynamic NeRF approaches Pumarola et al. (2021); Park et al. (2021a); Gao et al. (2021); Park et al. (2021b); Fridovich-Keil et al. (2023); Fang et al. (2022b); Cao and Johnson (2023) can model light transport but suffer from prohibitively slow rendering speed, along with inferior reconstruction quality compared to Gaussian Splatting methods Kerbl et al. (2023).
To address these limitations, we propose 4D Gaussian Ray Tracing (4D-GRT), a novel two-stage data generation pipeline that combines the reconstruction capabilities of Gaussian Splatting with physically-based ray tracing for camera effect simulation. To the best of our knowledge, 4D-GRT is the first work to perform ray tracing on dynamic scenes using a Gaussian scene representation. In the first stage, given multi-view video captured by pinhole cameras, we optimize a 4D Gaussian Splatting (4D-GS) representation using differentiable ray tracing. In the second stage, given camera effect parameters, we use ray tracing to simulate different camera effects on the reconstructed scene. This approach enables controllable and physically grounded camera effects generation, while achieving the fastest generation speed and better or comparable generation quality compared to baselines.
Our contributions are summarized as follows:
  • Problem Identification: Current video generation approaches are limited in understanding or generating videos with accurate camera effects given numerical parameters.
  • 4D-GRT Pipeline: We propose a novel data generation pipeline that integrates 4D-GS with ray tracing to rapidly generate high-quality videos with controllable camera effects.
  • Benchmark Dataset: We construct and release a comprehensive paired dataset containing 8 dynamic indoor scenes with multi-view videos under four camera effects.
2 Related Work
2.1 Generating Visual Data with Camera Effect
The generation of visual data with realistic camera effects has become increasingly crucial for training robust computer vision models. Unlike ideal pinhole camera models commonly assumed in computer vision, real cameras exhibit various optical phenomena causing different camera effects.
Synthetic Data Generation Approaches. Traditional approaches for generating camera effect data rely heavily on synthetic rendering pipelines using computer graphics software such as Blender Blender Online Community (2025). These methods provide precise control over camera parameters, generating numerically labeled ground truth data. However, such methods face several fundamental limitations: (1) high labor costs for creating realistic scenes, (2) a sim-to-real gap that limits model generalization.
Generative Model Approaches. The emergence of powerful generative models, particularly diffusion-based methods Rombach et al. (2022); Croitoru et al. (2023); Blattmann et al. (2023); Xing et al. (2024); Ho et al. (2022a); Chang et al. (2025); Ho et al. (2022b); Chu et al. (2023); Yeh et al. (2024), has opened new avenues for camera effect synthesis Fang et al. (2024). These approaches can generate diverse visual content with various camera effects. Recent representative works include Curved Diffusion Voynov et al. (2024), AKiRa Wang et al. (2025a), which enable control over optical distortions or camera effect parameters. However, generative methods suffer from several critical limitations for training data generation: (1) most focus on specific camera effects rather than general effect synthesis, (2) lack of precise physical constraints leading to unrealistic camera effects, (3) inability to accurately control camera parameters, and (4) difficulty in ensuring multi-view consistency or temporal consistency in dynamic scenes.
The above limitations motivate the need for approaches that can generate physically accurate visual data with camera effects in dynamic scenes. Our method is designed to address these challenges.
2.2 Dynamic Neural Rendering
Neural rendering has revolutionized scene reconstruction. This progression provides the necessary foundation for physically-based camera effect simulation.
NeRF. Neural Radiance Fields (NeRF) Mildenhall et al. (2021) introduced a new paradigm for representing static scenes, modeling them as continuous volumetric functions with MLPs that map 3D coordinates and viewing directions to density and color for photorealistic view synthesis. Numerous extensions Barron et al. (2021); Zhang et al. (2020); Yu et al. (2021); Müller et al. (2022); Sun et al. (2022); Barron et al. (2022); Su et al. (2024); Lin et al. (2025) address NeRF’s limitations, including adaptations for dynamic scenes Pumarola et al. (2021); Park et al. (2021a, b); Gao et al. (2021); Fang et al. (2022a); Liu et al. (2023); Yan et al. (2023); Song et al. (2023); Gan et al. (2023); Cao and Johnson (2023); Wang et al. (2023a); Chen et al. (2024). Among these, K-Planes Fridovich-Keil et al. (2023), HexPlane Cao and Johnson (2023), and MSTH Wang et al. (2023a) achieve state-of-the-art performance by leveraging plane- or grid-based representations to accelerate dynamic scene reconstruction while maintaining high rendering quality.
Gaussian Splatting. Despite NeRF’s impressive results, its high computational cost and implicit representation limit practical use. 3D Gaussian Splatting (3D-GS)Kerbl et al. (2023) addresses these issues with an explicit 3D Gaussian Ke et al. (2025) representation and efficient rasterization, enabling real-time, high-quality rendering Hou et al. (2025); Zhan et al. (2025). Recent worksYang et al. (2024b); Wu et al. (2024b); Liang et al. (2025); Huang et al. (2023); Lin et al. (2024); Kratimenos et al. (2024); Yang et al. (2024a); Fan et al. (2024) extend 3D-GS to dynamic scenes. For example, Deformable 3D Gaussians Yang et al. (2024b) introduce deformation modeling, while 4D Gaussian Splatting (4D-GS) Wu et al. (2024b) employs a spatial-temporal encoder composed of multi-resolution voxel planes and MLPs to model the deformation of 3D Gaussians. These advances preserve the efficiency of Gaussian Splatting while enabling high-quality dynamic scene reconstruction, establishing a strong foundation for camera effect simulation.
2.3 Camera effect rendering in neural scene representations
Traditional renderers simulate camera effects through ray tracing Cook et al. (1984), reproducing physical optics. In neural scene representations, NeRF can achieve similar effects Wu et al. (2022); Niu et al. (2024) due to its ray-casting–based rendering, but NeRF-based methods are computationally expensive and lack efficient rendering. In contrast, 3D-GS offers fast rasterization but is restricted to the pinhole camera model, requiring specialized designs and complex Jacobian computations to support effects such as fisheye distortion Liao et al. (2024), rolling shutter with motion blur Seiskari et al. (2024), and depth of field Lee et al. (2025); Wang et al. (2025b); Shen et al. (2025). 3DGRT Moenne-Loccoz et al. (2024b) overcomes this limitation by enabling direct ray tracing over 3D Gaussians, allowing physically accurate and flexible camera effect simulation, while 3DGUT Wu et al. (2025) achieves similar capabilities by approximating 3D Gaussians with sigma points via the unscented transform and projecting them through different camera models. However, both are restricted to static scenes and cannot simulate camera effects in dynamic settings. To address this, our method integrates 4D-GS with 3DGRT, leveraging ray tracing to enable camera effect rendering in dynamic scenes.
3 Method
Figure 3: The overall pipeline of our model. Given multi-view videos, we optimize the 4D-GS representation through differentiable ray tracing. Then, given camera effect parameters, we can utilize ray tracing to render videos with physically-correct camera effects.
As illustrated in Fig.3, our proposed 4D Gaussian Ray Tracing (4D-GRT) framework generates physically correct training data in two stages: dynamic scene reconstruction (Section 3.2) and camera-effect rendering (Section 3.3). In the first stage, we reconstruct a dynamic scene from synchronized multi-view videos using a 4D-GS Wu et al. (2024a) model with differentiable ray tracing Moenne-Loccoz et al. (2024a). In the second stage, we render the scene with physically based camera models, including fisheye distortion, depth of field, and rolling shutter, producing physically accurate videos for effect-aware vision training.
3.1 Preliminary
3D Gaussian Splatting (3D-GS). 3D-GS Kerbl et al. (2023) represents a static 3D scene as a collection of 3D Gaussians, each characterized by its mean position $(x,y,z)$, a covariance matrix $\mathbf{\Sigma}$ decomposed into scaling $s$ and rotation $r$ parameters, opacity $\sigma$, and spherical harmonic coefficients $\mathcal{C}$. Rendering is done by differentiable splatting Yifan et al. (2019), where $\mathbf{\Sigma}$ is projected into 2D as $\mathbf{\Sigma}^{\prime}=\mathbf{J}\mathbf{W}\mathbf{\Sigma}\mathbf{W}^{T}\mathbf{J}^{T}$, with $\mathbf{J}$ and $\mathbf{W}$ denoting the Jacobian of the projective transformation and the viewing transformation, respectively. Pixel colors are then computed via alpha compositing of the projected Gaussians from front to back, $C=\sum_{i\in\mathcal{N}}T_{i}\alpha_{i}c_{i}$, where $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance, $\alpha_{i}=\sigma_{i}\exp(-\tfrac{1}{2}\bm{x}^{T}\mathbf{\Sigma}^{\prime}\bm{x})$ is the opacity, and $c_{i}$ is the view-dependent color derived from spherical harmonics.
3.2 Dynamic Scene Reconstruction
As shown in Fig.  3, we reconstruct scene dynamics from multi-view videos $\{\hat{C}_{v,t}\mid v\in\mathcal{V},t\in\mathcal{T}\}$, where $\mathcal{V}$ denotes the set of camera views and $\mathcal{T}$ the set of time indices. Following 4D-GS Wu et al. (2024a), we represent the scene with a set of 3D Gaussians $G$ and a Gaussian deformation field network, which consists of a spatio-temporal structure encoder and a multi-head Gaussian deformation decoder to model the per-frame deformations of 3D Gaussians. Specifically, to compute the deformation of 3D Gaussians $G$ at time $t$, we first employ a 4D voxel planes $R_{l}$ to extract voxel features based on each Gaussian’s mean position and time step $t$. These features are first fused by a lightweight MLP $\varphi$, and then passed to separate MLPs $(\varphi_{x},\varphi_{r},\varphi_{s})$ that predict the residuals of the Gaussian attributes $((\Delta x,\Delta y,\Delta z),\Delta r,\Delta s)$. Finally, the attribute residuals are applied to $G$ to obtain the deformed Gaussians $G_{t}$. For details of the Gaussian deformation field network, please refer to the supplementary material.
For rendering, although 3D-GS Kerbl et al. (2023) rasterization provides efficient and high-quality results, it is inherently limited in its ability to simulate physically accurate camera effects. In particular, phenomena that require explicit ray-based modeling, such as complex lens distortions or light-transport-dependent optical effects, cannot be faithfully reproduced. To address these limitations, we adopt the differentiable ray tracing framework of 3DGRT Moenne-Loccoz et al. (2024a), which traces rays directly through 3D Gaussian primitives instead of projecting them onto the image plane. By combining a $k$-buffer hit-based marching scheme with the hardware-accelerated NVIDIA OptiX interface, this approach achieves both the accuracy required for realistic camera-effect rendering and the efficiency necessary for large-scale data generation.
By integrating the dynamic scene model with differentiable ray tracing, we reconstruct dynamic scenes end-to-end via training with a color loss. At each training iteration, we randomly sample a view $v\in\mathcal{V}$ and a time $t\in\mathcal{T}$, obtain the deformed Gaussians $G_{t}$, and render them via ray tracing from view $v$ to obtain the image $C_{v,t}$. We then compute the L1 loss between $C_{v,t}$ and the ground-truth image $\hat{C}_{v,t}$, denoted by $\mathcal{L}_{1}$. Additionally, we apply a grid-based total variation loss $\mathcal{L}_{\mathrm{TV}}$ following Wu et al. (2024a); Cao and Johnson (2023). The overall loss function is defined as
$\mathcal{L}=\mathcal{L}_{1}(C_{v,t},\hat{C}_{v,t})+\mathcal{L}_{\mathrm{TV}}.$ (1)
3.3 Camera Effects Rendering
After reconstructing the dynamic scene, we synthesize common camera effects by integrating physical camera models into a ray-tracing renderer. Specifically, we implement three representative effects: fisheye distortion, depth of field, and rolling shutter.
Fisheye. Fisheye lenses enable extremely wide-angle capture but introduce strong radial distortion. To simulate this effect, we follow Blender Blender Documentation Team (2025) and adopt a 4th-degree polynomial radial distortion model, which provides a flexible and physically grounded representation of fisheye lens characteristics. The model is defined as:
$\theta$ $=k_{0}+k_{1}r+k_{2}r^{2}+k_{3}r^{3}+k_{4}r^{4},$ (2)
where $r$ is the radial distance from the principal point of the camera sensor, $\theta$ represents the polar angle between the optical axis and the direction of the incoming ray, and $k_{i}$ are the distortion polynomial coefficients that define the specific fisheye lens characteristics.
Given the sensor dimensions and coefficients $k_{i}$, we convert each pixel’s image coordinates into physical sensor coordinates $(x,y)$ (in millimeters). We then compute the radial distance $r=\sqrt{x^{2}+y^{2}}$, evaluate the polynomial to obtain the polar angle $\theta$, and calculate the azimuthal angle $\phi=\arccos(x/r)$. These angles define the spherical ray direction, which is then passed to the ray tracer for rendering. Compared to prior generative methods, this parameter-based approach follows real lens behavior and preserves consistency across multiple views.
Depth of Field (DoF). DoF controls which parts of a scene appear sharp and which appear blurred, allowing cameras to emphasize subjects at specific depths while de-emphasizing background or foreground regions. We simulate this effect using the model introduced in Cook et al. (1984). For each pixel, given the focus distance $f_{z}$ and the aperture radius $r_{a}$, we first compute the intersection point $\mathbf{p}=\mathbf{o}+f_{z}\mathbf{d}$ of the ideal pinhole ray with the focal plane at distance $f_{z}$. We then jitter the ray origin over a circular aperture of radius $r_{a}$ by sampling a point $\bm{\ell}=(\ell_{x},\ell_{y})$ uniformly from the unit disk and mapping it to world space using the camera’s lens-plane basis $(\mathbf{u},\mathbf{v})$, where $\mathbf{u}$ and $\mathbf{v}$ are orthogonal unit vectors spanning the lens plane. The perturbed ray origin and direction are then computed as
$\mathbf{o}^{\prime}=\mathbf{o}+\ell_{x}\mathbf{u}+\ell_{y}\mathbf{v},\qquad\mathbf{d}^{\prime}=\frac{\mathbf{p}-\mathbf{o}^{\prime}}{\lVert\mathbf{p}-\mathbf{o}^{\prime}\rVert}.$ (3)
Repeating this process for multiple samples per pixel and averaging their traced radiance yields physically accurate defocus blur, with the blur size determined by the aperture radius and the object’s depth relative to the focal plane.
Rolling Shutter. The rolling shutter effect arises because image rows are not captured simultaneously but sequentially over time. As a result, objects in motion or a moving camera can cause geometric distortions in the recorded image. To simulate this effect, we cast a single ray for each pixel through the deformed 3D Gaussians at its corresponding capture time. This approach resembles the motion blur technique of Cook et al. (1984), but it differs in that we assume no motion blur: instead of integrating multiple rays over the shutter interval of a row, we trace only one ray per pixel at the corresponding sensing time. Concretely, for a pixel at row $r$, we first compute the sensing time $t_{r}$, then evaluate the deformed 3D Gaussians $G_{t_{r}}$. A ray is traced through the deformed 3D Gaussians $G_{t_{r}}$ to determine the pixel’s radiance. We currently assume a static camera; however, the method can be extended to a moving camera by interpolating the camera trajectory at time $t_{r}$ and casting the ray from the position.
A limitation of this approach is that sensing times differ across rows, preventing full parallelization of ray tracing and thus limiting rendering speed. To address this, we adopt an approximation strategy. Specifically, we divide the rows into chunks of size $N_{c}$. Assuming the scene motion is moderate and the shutter time is relatively short, we approximate the sensing time of all rows in a chunk by their average. For each chunk, we first compute the deformed Gaussian at this average sensing time and then trace the rays within the chunk in parallel. This strategy greatly improves rendering efficiency, though an excessively large chunk size may introduce blocking artifacts due to the approximation.
4 Experiments
Datasets. In real-world scenarios, simultaneously capturing dynamic scenes with different camera effects is extremely challenging. Existing public datasets lack multi-view data for the same dynamic scene under different camera effects. To address this limitation, we constructed our own dataset using Blender 4.5 Blender Online Community (2025), which enabled us to: (1) simulate realistic physics, and (2) consistently reproduce identical dynamic scenes while applying various camera effects.
We constructed 8 distinct dynamic scenes across 4 indoor environments (basketball court, warehouse, living room and bathroom) using high-quality models from BlenderKit BlenderKit Community (2025) and other asset platforms. Each scene features diverse material properties and physical interactions, including bouncing balls, mechanical animations, and wind-influenced motions. For each scene, we established 50 camera viewpoints rendered under four different camera effects (pinhole, fisheye, rolling shutter, depth of field), generating 50-frame videos at 512×512 resolution. Camera placement and point cloud generation were automated using the PlenoBlenderNeRF plugin H. (2025); Raafat (2024), ensuring consistent geometric foundations across different camera effects. Detailed scene descriptions and technical implementation are provided in the supplementary material.
Table 1: Quantitative comparison on pinhole camera rendering.
[…]
HexPlane Cao and Johnson (2023) 23.1124 0.7956 0.2942 0.20
MSTH Wang et al. (2023a) 29.431 0.9023 0.1139 9.38
4D-GRT (Ours) 32.801 0.8898 0.1018 36.56
Table 2: Quantitative comparison on depth-of-field effect rendering.
[…]
HexPlane Cao and Johnson (2023) 18.3722 0.7343 0.5056 0.01
MSTH Wang et al. (2023a) 28.4692 0.9009 0.1540 0.57
4D-GRT (Ours) 31.2475 0.9124 0.1210 3.44
Table 3: Quantitative comparison of fisheye distortion rendering. Metrics with "_m" indicate that the metric is computed within the pre-defined mask.
[…]
HexPlane Cao and Johnson (2023) 15.6489 0.6678 0.4142 21.6869 0.7480 0.2726 0.21
MSTH Wang et al. (2023a) 24.4222 0.8163 0.1764 26.7930 0.8571 0.1153 9.38
4D-GRT (Ours) 24.1322 0.8162 0.1678 28.8927 0.8555 0.1259 41.53
Table 4: Quantitative comparison on rolling shutter effect.
[…]
HexPlane Cao and Johnson (2023) N/A 21.3468 0.7723 0.3162 0.21
MSTH Wang et al. (2023a) N/A 28.7003 0.8927 0.1190 9.35
4D-GRT (Ours) 1 row 31.6144 0.8821 0.1056 0.76
4D-GRT (Ours) 4 rows 31.6146 0.8821 0.1056 4.99
4D-GRT (Ours) 16 rows 31.6082 0.8820 0.1056 13.54
Table 5: Comparison of training time and storage.
[…]
HexPlane Cao and Johnson (2023) 12 hours 238
MSTH Wang et al. (2023a) 8 mins 889
4D-GRT (Ours) 3 hours 364.875
Baselines. Our task of generating videos with camera effects from dynamic scenes is novel. For fair and meaningful evaluation, we compare against two state-of-the-art dynamic NeRF methods, HexPlane Cao and Johnson (2023) and MSTH Wang et al. (2023a), selected for their capabilities in 4D scene reconstruction and ray tracing. To ensure consistency, we apply the same camera effect rendering module used in our method to these baselines, eliminating differences in effect generation. We conduct both quantitative and qualitative comparisons.
Evaluation Protocol. All methods are trained on multi-view videos rendered with a pinhole camera on GeForce RTX 4090. For evaluation, we render the same camera viewpoints with different camera effects (fisheye, rolling shutter, and depth of field) on GeForce RTX 4090, and then compute rendering FPS and the metrics against the corresponding ground-truth videos. For both training and evaluation, we use all camera viewpoints in all time frames. To ensure fairness, all methods, including ours, are trained on the identical set of input videos until convergence.
Evaluation Metrics. We employ a set of metrics to assess the visual quality of our rendered videos against the ground-truth data. Specifically, we use Peak Signal-to-Noise Ratio (PSNR) Huynh-Thu and Ghanbari (2008), Structural Similarity Index Measure (SSIM) Wang et al. (2004), and Learned Perceptual Image Patch Similarity (LPIPS) Zhang et al. (2018). We also report training time, inference speed (FPS), and storage for efficiency evaluation. For all metrics, we report the average score over all frames, camera views, and scenes. For the rolling-shutter rendering comparison, we evaluate our method with chunk sizes of $N_{c}=1$, $4$, and $16$.
Quantitative comparison. Tables 2, 2, 3, 5, and 5 present the quantitative results. Across all camera effects including pinhole, fisheye, depth-of-field and rolling-shutter rendering, our method delivers the highest rendering speed. Besides this advantage, the rendering quality of ours remains competitive. For pinhole, depth-of-field, and rolling-shutter rendering, 4D-GRT achieves higher PSNR than both baselines, with SSIM and LPIPS comparable to MSTH. For fisheye rendering, it matches MSTH across all metrics while still outperforming HexPlane. The rendering speedup stems from our hardware-accelerated Gaussian ray tracing, while the quality gains follow from the continuous 4D-Gaussian representation.
In fisheye evaluation, the larger field of view reveals regions not visible in the pinhole cameras used for training. Since these regions are never observed, their appearance is unconstrained and leads to unreliable metrics. To address this, we adopt a masked evaluation strategy. Specifically, a circular mask centered at the principal point is applied to both the rendered and ground-truth frames, restricting the comparison to regions visible to the training cameras. We empirically set the diameter of the circle as 409.6. In this evaluation, as shown in Table 3, our method achieves substantially higher PSNR than the baselines, with comparable SSIM and LPIPS.
Regarding rolling-shutter rendering FPS, our method without approximation is slower than the baselines due to structural differences. Specifically, our approach models the dynamic scene using a set of canonical 3D Gaussians and a Gaussian deformation field network, which requires generating a separate deformed scene for each row in the rolling-shutter image, making parallel ray tracing infeasible. In contrast, the baselines employ plane-based or grid-based 4D representations, allowing direct spatio-temporal queries per ray, benefiting from parallelism for faster rendering. However, as shown in Table 5, we achieve higher FPS with only minimal quality loss with our approximation strategy. Larger chunk sizes yield higher FPS at the cost of slight image quality degradation, while smaller chunk sizes preserve higher rendering quality at the expense of longer rendering times.
In terms of efficiency, HexPlane is the most storage-efficient but slowest to train, while MSTH trains fastest but requires the most storage. Our method achieves the highest reconstruction and rendering quality with fast rendering speed, while keeping both training time and storage at reasonable levels.
Pinehole Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rolling
Shutter
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Depth
of
Field
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fisheye Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth Hexplane Cao and Johnson (2023) MSTH Wang et al. (2023a) 4D-GRT(Ours) Ground Truth Hexplane Cao and Johnson (2023) MSTH Wang et al. (2023a) 4D-GRT(Ours)
Pinehole Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Rolling
Shutter
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Depth
of
Field
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fisheye Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative comparison on our synthetic datasets. The artifacts are framed by red rectangles
Qualitative comparison. For qualitative evaluation, we present the rendering results of 4D-GRT (ours), MSTH Wang et al. (2023a), and Hexplane Cao and Johnson (2023) in Fig. 4. Our method demonstrates higher-fidelity results compared to others. Specifically, the reconstruction quality of Hexplane Cao and Johnson (2023) is obviously worse than ours, and there are reconstruction artifacts and color temperature misalignment in the results of MSTH Wang et al. (2023a), which do not exist in our results.
Qualitative results on real-world dataset. We further validate on real-world dynamic scenes by reconstructing and rendering sequences from the Neural 3D Video (Plenoptic Video) dataset Li et al. (2022) with camera effects in Fig. 5; more qualitative examples are provided in the supplementary material. In addition, we present results as videos to better demonstrate temporal consistency, and also include rendering results on sequences from our own dataset in video form for visualization of scene dynamics.
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Pinhole Depth of Field Rolling Shutter Fisheye
Figure 5: Qualitative results of 4D-GRT on the Neural 3D Video dataset  Li et al. (2022).
5 Conclusion and Limitations
Conclusion. In this work, we identify a critical bottleneck: contemporary world models, built on pretrained video generators, fail to respect camera-effect parameters and thus produce physically inconsistent videos. To address this, we introduced 4D Gaussian Ray Tracing (4D-GRT), a two-stage pipeline that couples 4D Gaussian splatting with differentiable ray tracing for controllable, effect-accurate data generation. We also release a benchmark of 8 dynamic indoor scenes with synchronized multi-view videos rendered under four camera effects. Experimentally, 4D-GRT delivers substantially higher rendering speed than strong dynamic NeRF baselines across four types of camera effect while maintaining competitive or better image quality (PSNR/SSIM/LPIPS). This advancement demonstrates that our method is more suitable for generating substantial amounts of videos with camera effects, which can better address the data insufficiency problem in understanding camera effects across world models and other visual systems. Together, 4D-GRT and our benchmark provide a strong foundation for advancing camera-aware vision.
Limitations. Our method requires well-reconstructed dynamic scenes to render high-quality videos with camera effects, necessitating sufficient multi-view video data. This dependency limits its applicability in scenarios with sparse viewpoints or monocular inputs. Future work could alleviate this limitation by integrating generative or foundation models to reconstruct dynamic scenes from monocular video, making camera effect simulation more accessible across diverse scenarios.
Supplementary Material Overview
This supplementary document is structured as follows:
  1. A.
    Dataset Construction
  2. B.
    Video Generation Experiment
  3. C.
    Implementation Detail
  4. D.
    Additional Qualitative Results
Appendix A Dataset Construction
In practice, capturing the same dynamic scene with different camera effects is infeasible; re-takes change motion timing, illumination, and other conditions. Public datasets therefore lack multi-view, same-scene recordings under multiple effects.
To fill this gap, we construct a synthetic multi-view dataset in Blender 4.5 Blender Online Community (2025). We create 8 dynamic indoor scenes across 4 environments (basketball court, warehouse, living room, bathroom), each with 50 viewpoints and 50 frames at $\times512$, rendered under four effects (normal, fisheye, rolling shutter, depth of field).
Section A.1 details the Blender Blender Online Community (2025) pipeline, tools, and plugins used for multi-view dynamic rendering, and Section A.2 describes the scenes and their material/physics characteristics.
A.1 Scene Construction using Blender
To establish a dataset suitable for evaluating model performance under different camera effects, we selected Blender 4.5 Blender Online Community (2025) as our primary scene construction and rendering tool. Blender provides highly realistic physics simulation capabilities, enabling accurate modeling of object motion, collisions, bouncing, and other dynamic behaviors. Additionally, it ensures completely consistent scene configurations across multiple renders, guaranteeing perfect content alignment under different camera effects while avoiding the limitations of real-world filming where identical dynamics cannot be reproduced.
For scene sourcing, we selected 4 high-quality indoor scene models through the BlenderKit BlenderKit Community (2025) asset library. In addition, certain dynamic objects, such as the basketball model used in the basketball court scene, were also obtained from BlenderKit. In the scenes of living room and bathroom, the dynamic objects were obtained from other online asset platforms: the LEGO bulldozer model Heinzelnisse (2014) and the Box animation model RonKo (2023) from Blend Swap BlendSwap Community , the running cat model dailyfree3d (2019) and a pot plant model 3dhaupt (2019) from free3D.com Free3D . Within each indoor environment, we designed 2 different types of dynamic objects and animations, such as bouncing spheres, jumping cubes, or other wind-influenced motion behaviors. In total, we constructed 8 distinct dynamic scenes.
We chose 8 scenes primarily because they encompass diverse material properties and physical interaction patterns, including reflective floor surfaces and rigid body collision dynamics. This diversity adequately simulates various common lighting and motion scenarios found in the real world, providing comprehensive coverage for subsequent model generalization testing. Simultaneously, maintaining a reasonable number of scenes ensures that each scene’s dynamic design and camera configuration can undergo thorough manual adjustment and verification, preventing quality degradation due to excessive scene quantities.
For camera configuration, we established 50 different camera viewpoints for each scene, with each viewpoint rendered separately under four camera effects: pinhole, fisheye, rolling shutter, and depth of field. We selected 50 viewpoints to maintain manageable data volumes while covering sufficient observation angles, ensuring comprehensive capture of dynamic object shape variations and lighting responses. This design also provides adequate observational diversity for our 4D Gaussian ray tracing model, facilitating learning of complete 3D-4D scene structure during training.
Our 50 camera viewpoint placement, corresponding camera pose data, and initial scene point cloud generation were all accomplished through the PlenoBlenderNeRF H. (2025) Blender plugin that is originally forked from BlenderNeRF Raafat (2024). This plugin automatically distributes multiple cameras uniformly throughout the scene and generates corresponding initial point clouds from Blender mesh vertex data before rendering, ensuring complete consistency between rendering results under different camera effects and geometric foundations, facilitating subsequent quality analysis and comparison.
For data output, we generated 50-frame videos at 512×512 resolution for each camera under all four camera effects in each scene, ensuring complete capture of dynamic object motion trajectories, lighting changes, and camera effect characteristics for subsequent analysis.
A.2 Scene Descriptions and Characteristics
Our 8 constructed dynamic scenes derive from 4 different environment types: basketball court, warehouse, living room, and bathroom. Each scene incorporates distinct physical interaction patterns and surface properties in modeling and material settings, ensuring dataset coverage of diverse lighting responses and motion behaviors. The following sections describe each scene’s dynamic design and attributes.
Basketball Court:
  1. 1.
    Rolling Basketball: We simulate external forces acting on the basketball surface using Blender’s wind force, generating rolling motion on the ground. During rolling, the ball experiences both friction and inertia effects, creating gradually decelerating dynamic effects that closely approximate real physical behavior.
  2. 2.
    Falling Basketballs: The scene contains three basketballs positioned at height, falling under gravity to the ground surface. After elastic collisions with the floor, the basketballs interact through further collisions, causing directional changes and eventual rolling in different directions. This design simultaneously simulates multiple physical phenomena including multi-body collisions, bouncing, and rolling, effectively testing model capabilities in handling multi-target dynamics and interactions during rendering and geometric reconstruction.
Warehouse:
  1. 1.
    Bouncing Spheres: Three rigid spheres positioned at height fall under gravity to the ground surface, undergoing elastic collisions followed by mutual interactions. The collision process alters sphere motion directions and velocities, causing final rolling in different directions and creating complex multi-body interaction and scattering dynamics.
  2. 2.
    Bouncing Cubes: Four cubes fall from height to the ground surface, experiencing elastic collisions and mutual interactions. During collisions, cube motion trajectories and rotation angles continuously change, simulating non-spherical rigid body motion characteristics under multiple collision scenarios.
Living Room:
  1. 1.
    LEGO Bulldozer Model: We utilize a LEGO bulldozer model as the primary dynamic object in the living room scene. Through animation design, we simulate mechanical arm swinging motions to present realistic mechanical operation dynamics and joint motion trajectories while preserving material details and lighting variations in rendering.
  2. 2.
    Running Cat Model: We also employ a cat model with added temporal displacement and running animations, creating continuous motion postures within the scene. The animation encompasses both overall positional changes and continuous limb movements with posture variations, simulating realistic animal running physics and rhythmic characteristics.
Bathroom:
  1. 1.
    Rotating Pot Plant Model: We utilize a detailed pot plant model as an additional dynamic element in the bathroom scene. Using Blender’s key frame animation system, we create a smooth rotational motion where the pot plant rotates 180 degrees around the z-axis throughout the animation sequence. This rotation provides a controlled, predictable motion pattern that allows for systematic evaluation of how different camera effects handle circular motion and changing viewpoints of complex organic geometry.
  2. 2.
    Box animation Model: The second dynamic element in the bathroom scene is a self-opening box whose lid gradually reveals its interior. The motion involves coordinated hinge movement, producing geometric changes, occlusion effects, and shifting light as inner surfaces are exposed—ideal for testing camera effects on mechanical motion and surface transitions.
Appendix B Video Generation Experiment
To evaluate the capabilities of state-of-the-art video generation models in synthesizing videos with realistic camera effects, we conduct comprehensive experiments using four leading commercial video generation models: OpenAI Sora OpenAI (2024), Google Veo 3 Google DeepMind (2025), Tencent HunyuanVideo Kong et al. (2025), and ByteDance Seedance 1.0 Mini Gao et al. (2025). Our evaluation focuses on three fundamental camera distortion effects: fisheye distortion, rolling shutter artifacts, and depth of field effects. These effects represent critical aspects of realistic video synthesis that require accurate understanding of camera optics and sensor behavior.
B.1 Video Generation Configuration
To ensure fair comparison while respecting each model’s individual design constraints and optimal operating conditions, we utilize the default output specifications for each video generation model. This approach allows us to evaluate the models under their native configurations rather than forcing uniform parameters that may disadvantage certain architectures.
B.2 Experimental Design and Prompting Strategy
We design a systematic evaluation protocol using three distinct prompts for each camera effect to assess both implicit understanding and explicit parameter-driven generation capabilities. This multi-prompt approach enables us to evaluate model performance across different levels of technical specification complexity.
Figure 6: Fisheye video evaluation. Sora most closely follows the specified lens parameters, while the others often fail to produce a true fisheye effect. Each cell shows the 60th frame of the generated video.
B.2.1 Fisheye Distortion Generation
To evaluate fisheye rendering, we use three prompts of increasing specificity. Figure 6 shows the video results; we use the 60th frame from each video as the representative image.
Basic prompt.
We first probe implicit fisheye understanding with a natural-language description:
“A fisheye-lens video of a person skateboarding in a city square.”
Detailed camera specification (extreme wide-angle).
We test explicit parameter control using an extreme wide-angle setup:
“A video of a person skateboarding in a city square, captured using an **8mm fisheye lens** with a **180° field of view**.”
Detailed camera specification (moderate wide-angle).
We also assess sensitivity with a less extreme configuration:
“A video of a person skateboarding in a city square, captured using a **15mm fisheye lens** with a **150° field of view**.”
Findings.
Across the three prompts, Sora is the only model that consistently produces recognizable fisheye imagery and approximately respects the specified parameters. The other models (Veo 3, HunyuanVideo, and Seedance 1.0 Mini) frequently fail to render a convincing fisheye effect or to follow the camera settings. In particular, HunyuanVideo under the basic prompt shows no fisheye distortion, and Seedance 1.0 Mini under the moderate wide-angle prompt yields only a generic wide-angle view, losing the characteristic fisheye curvature (see Fig. 6).
B.2.2 Rolling-Shutter Effect Generation
For rolling-shutter evaluation, we use three prompts that vary only the strength of the effect. Figure 7 shows the results from four models (3 prompts × 4 models = 12 images); for visualization, we display the 60th frame from each generated video.
Basic prompt.
We probe implicit rolling-shutter understanding with a motion-centric description:
“A rolling-shutter video of an electric fan spinning at high speed.”
Detailed camera specification (pronounced effect).
We enforce stronger distortion by specifying a longer readout:
“A rolling-shutter video of an electric fan, captured with a **readout time of 40 ms**.”
Detailed camera specification (subtle effect).
We reduce the distortion by shortening the readout:
“A rolling-shutter video of an electric fan, captured with a **readout time of 10 ms**.”
Findings.
Across the three prompts, only Veo 3 reliably exhibits the characteristic rolling-shutter skew in the fan blades. The other models (Sora, HunyuanVideo, Seedance 1.0 Mini) produce blades with uneven apparent areas and lack the expected geometric skew, indicating poor adherence to the rolling-shutter effect (see Fig. 7).
Figure 7: Rolling-shutter video evaluations. Only Veo 3 exhibits the expected rolling-shutter skew in the fan blades; the other models produce uneven blade areas without the characteristic skew.
B.2.3 Depth-of-Field Effect Generation
For depth-of-field evaluation, we use three prompts that test models’ understanding of blur and focus control. Figure 8 shows the results from the four models; for visualization, we display the 60th frame from each generated video.
Basic prompt.
We first probe implicit depth-of-field understanding with a natural description:
“A video of a chess piece on a board with very shallow depth of field.”
Detailed camera specification (shallow focus).
We test explicit parameter control for shallow depth of field:
“A video of a chess piece on a board, shot with a **35 mm focal length** and **aperture f/2.0**.”
Detailed camera specification (moderate focus).
We evaluate sensitivity under a smaller aperture:
“A video of a chess piece on a board, shot with a **35 mm focal length** and **aperture f/8.0**.”
Findings.
Across all three prompts—including the explicit f/2.0 (shallow) and f/8.0 (moderate) settings—every model renders pronounced background blur in the 60th-frame images, regardless of the specified aperture. This indicates that the models do not respect aperture parameters or depth relationships, and therefore perform poorly on depth-of-field video generation (see Fig. 3).
In summary, across fisheye, rolling-shutter, and depth-of-field evaluations, current video generation models exhibit limited camera awareness. Only Sora reliably produces recognizable fisheye imagery, and only Veo 3 consistently displays the expected rolling-shutter skew; depth-of-field outputs remain strongly blurred regardless of the specified aperture (e.g., ff/2.0 vs. ff/8.0). These results indicate poor compliance with camera parameters and insufficient physical fidelity, underscoring the need for camera-aware training data and modeling approaches.
Figure 8: Depth-of-field video evaluation. All models produce pronounced blur regardless of the specified aperture (ff/2.0 vs. ff/8.0), revealing weak compliance with DoF parameters and poor control of depth-of-field effects.
Appendix C Implementation Detail
C.1 Gaussian Deformation Network Architecture
We adopt the 4DGS Wu et al. (2024b) formulation of the Gaussian deformation network to predict the deformation of 3D Gaussians at a specific time step. Specifically, the network consists of a spatio-temporal structure encoder and a multi-head Gaussian deformation decoder.
The spatio-temporal encoder is constructed using six multi-resolution plane modules
$R_{l}(i,j)\in\mathbb{R}^{h\times lN_{i}\times lN_{j}},$
and an MLP $\varphi$, where $l$ is the upsampling scale, $h$ is the dimension of the voxel feature, and $N_{i}$ and $N_{j}$ are the base resolutions of the plane. Given the mean value $(x,y,z)$ of a 3D Gaussian
$G=\{(x,y,z),r,s,\sigma,\mathcal{C}\},$
we first compute the voxel feature
$f_{h}$ $=\bigcup_{l}\prod\mathbf{interp}(R_{l}(i,j))\in\mathbb{R}^{h\times l},$ (4)
$(i,j)$ $\in\{(x,y),(x,z),(y,z),(x,t),(y,t),(z,t)\},$ (5)
where $\mathbf{interp}$ denotes bilinear interpolation. The feature $f_{h}$ is then merged by $\varphi$ to produce
$f=\varphi(f_{h}).$
Next, the multi-head Gaussian deformation decoder predicts the deformation of 3D Gaussians. The decoder consists of separate MLPs $\{\varphi_{x},\varphi_{r},\varphi_{s}\}$, which predict the deformation residuals of each Gaussian attribute:
$(\Delta x,\Delta y,\Delta z)=\varphi_{x}(f),\quad\Delta r=\varphi_{r}(f),\quad\Delta s=\varphi_{s}(f).$
These residuals are then applied to deform a 3D Gaussian $G$ into
$G_{t}=\{(x+\Delta x,y+\Delta y,z+\Delta z),\,r+\Delta r,\,s+\Delta s,\,\sigma,\,\mathcal{C}\}.$
C.2 Training details
The hyperparameters in our experiments are mainly based on those reported in 4DGS Wu et al. (2024b). In our observation, we find the setting following 4DGS Wu et al. (2024b) will not perform well in scenes with rapidly moving objects. Therefore, we increase the learning rate, the capacity of the deformation field, the basic resolution of the plane modules and the training iterations. Specifically, we adopt the following setting for scenes with objects undergoing vigorous motion. The width of deformation MLP is set to $128$. The basic resolution of the plane modules is set to $128$, which is upsampled by $2$, 44 and $8$. The densification threshold is set as $2.5\times 10^{-5}$ and increases to $1\times 10^{-4}$. The training iteration in the coarse stage is set to $5\times 10^{4}$ and in the fine stage it is set to $8\times 10^{4}$.
We adopt a two-stage optimization strategy similar to 4DGS Wu et al. (2024b). In the first stage, we initialize the positions and colors of the 3D Gaussians using point clouds generated from Blender’s synthetic scenes. We then train the 3D Gaussian scene—without the deformation network—using only the first frame. This step reconstructs the static components of the scene and provides a strong initialization for dynamic objects. In the second stage, we introduce the deformation network and jointly optimize both the 3D Gaussians and the network across all time frames.
In 3DGS Kerbl et al. (2023), densification is guided by the magnitude of the 2D positional gradients of Gaussians. In contrast, since our rendering relies on ray tracing rather than rasterization, we follow the approach of Moenne-Loccoz et al. (2024a) and use the magnitude of the 3D gradient of the Gaussian means to decide when to densify a Gaussian. Additionally, as in Moenne-Loccoz et al. (2024a), we scale the gradient magnitude by the distance of the deformed Gaussian from the camera, encouraging densification in the background.
Appendix D Additional Qualitative Result
We present a qualitative comparison of rolling-shutter rendering with different chunk sizes. As shown in 9, smaller chunk sizes produce more accurate renderings, while larger chunk sizes lead to noticeable blocking artifacts.
Figure 9: Qualitative comparison of rolling shutter effect with different number of rows per chunk.